Let u =

Let u = < -3, -3> , v = < 9, -5> . Find 6u - 4v
Answer #1

Solution: given that U= -3,-37 IV = 29,-57 we find 64-4V-? NOW SONA 6u = 62-3,-37 = 2-18,5187 and 4V - 429,-57 = 236,-207 We